OpenStudy (anonymous):

what are the coordinates of the vertex function: y-11=-(x-5)^2

OpenStudy (dan815):

hey

OpenStudy (anonymous):

hi!

OpenStudy (dan815):

(x-5) means the graph is shifted to what direction?

OpenStudy (anonymous):

it is negative right? and to the left?

OpenStudy (dan815):

works a little hte other way around

OpenStudy (anonymous):

so the right?

OpenStudy (dan815):

yeah

OpenStudy (dan815):

think about it for x^2

OpenStudy (dan815):

if u put 0 in for y=x^2, the aswer is 0^2 = 0 but for (x-5)^2 ; when x = 0 ; y = 25!

OpenStudy (anonymous):

so its, (0,25) ?

OpenStudy (dan815):

and u dont get to 25 on y=x^2 untill you get to x = 5

OpenStudy (dan815):

yes good observation!

OpenStudy (dan815):

okay lets see how the other values effect this graph

OpenStudy (anonymous):

haha! thank you!

OpenStudy (dan815):

y-11=-(x-5)^2 theres a - infront of that whole (x-5)^2 so its a negative parabola function(going downward)

OpenStudy (anonymous):

oh okay that makes sense

OpenStudy (dan815):

and the value of y is affected by this y-11, how does this effect the graph

OpenStudy (anonymous):

also facing downward?

OpenStudy (dan815):

remember how (x-5) told you some shift to the right or left what about (y-11) will that also tell you some shift, perhaps up or down now

OpenStudy (anonymous):

down!

OpenStudy (dan815):

again its a little bit weird like the x-5, its actually being shifted up

OpenStudy (dan815):

but its a little more easier to see if you put it in this form

OpenStudy (dan815):